suppose you bought supplies for a party. 3 rolls of streamers and 15 party hats cost $30. later you bought 2 rolls of streamers and 4 party hats for $11. how much did each roll of streamers cost. how much did each party hat cost?
Answer by AnlytcPhil(1116)

suppose you bought supplies for a party. 3 rolls of streamers and 15 party hats
cost $30. later you bought 2 rolls of streamers and 4 party hats for $11. how
much did each roll of streamers cost. how much did each party hat cost?

>>...3 rolls of streamers and 15 party hats cost $30...<<

Translation into algebra:  3R + 15H = 30

>>...later you bought 2 rolls of streamers and 4 party hats 
for $11...<<

Translation into algebra:   2R + 4H = 11

3R + 15H = 30
2R +  4H = 11

Can you solve that system?  If not repost.

Answer R = 5/2 or 2 1/2 rolls
       H = 3/2 or 1 1/2 party hats

Something is obviously wrong.  While it is feasible to buy 
2 1/2 rolls of streamers, it is not feasible to by 1 1/2 
party hats.

Edwin
AnlytcPhil@aol.com